OpenStudy (lilai3):

PLEASEPLEASEPLEASE HELP!!!! will award medal/fan. Calvin invested 7500 for one year, part at 12% annual interest and the rest at 10% annual interest. His total interest for the year was $890. How much money did he invest at 12%? Thank you.~

OpenStudy (anonymous):

Let's say the amount he invested for 12% annual interest was x. Then, the amount he invested for 10% annual interest was 1-x. We have the equation: \[\frac{ 12 }{ 100 }\left( 7500x \right) + \frac{ 10 } {100}\left( 7500(1-x) \right) = 890\] Solve using simple algebra. He invested ~93.33..% at 12% interest. The rest at 10%.

OpenStudy (lilai3):

is it 0.22x+890=7500??

OpenStudy (rane):

yea @lilai3

OpenStudy (lilai3):

it is? i just guessed lol is @haziqmir right too though?

OpenStudy (rane):

well, your way is kind a easy but im not sure of u'll get full marks in the tests/ exams but @haziqmir done it step by step and it will be better if u use that method and yes its correct

OpenStudy (lilai3):

can you please restep?

OpenStudy (lilai3):

i don't really get his way tho lol . .

OpenStudy (rane):

y not ?

OpenStudy (lilai3):

it doesn't just....make sense.... please explain bcuz my test is 2mr, and i'm really worried cuz i don't get thsese sorta problems. pelase help

OpenStudy (rane):

i'll let @haziqmir do that, since he was already doing it, i dont want to take over his explanation @lilai3 r u gona do it @haziqmir

OpenStudy (rane):

he's here

OpenStudy (anonymous):

I did post a reply. Didn't get through I guess. I am going to simplify this. Calvin has $7500. He wants to invest it in two parts – one at 12%, one at 10%. We are interested in finding out how much we invested in for 12%. So, we'll let that amount be x. After investing x at 12%, he has some amount left, which is 1-x, that he invests at 10%. Both annually. At the end of the year, he has a total interest of $890 from both the amounts. We know the interest percentage. The total interest he received will be equal to the interest received from both the amounts, which is 12% of amount x of 7500 AND 10% of the remaining amount. Or, the equation I wrote above. You can solve this from here on using algebra. Am I clear now?

OpenStudy (lilai3):

ohhhhhhhhhhh i get it. is it $7,000?

OpenStudy (lilai3):

thank you very much. it's l8 in my side of the world. thanks 4 helping me. here's a medal.(; good nite

OpenStudy (anonymous):

Phew. :) Good luck for your exams. And thank YOU :)

OpenStudy (anonymous):

As for 1-x, think of it like this: The maximum amount you can take of $7500 is $7500 itself, which is 100%, or 1 times $7500. When you took x out of it, $7500 turns into 7500(1-x). Hope that didn't confuse you. It all makes sense.

OpenStudy (rane):

FINALLY! anyway good luck with your exams @lilai3

OpenStudy (lilai3):

thank you! thanks to all of you! >.<